OpenStudy (3psilon):

How can you turn sin5x/9x into sin(5x) / (9x) (5/9) sin(5x)/(5x) ???? I don't get how this works out mathematically

OpenStudy (zehanz):

Multiply with 9, multiply with 5:\[\frac{ \sin 5x }{ 9x }=\frac{9 \cdot 5\sin5x}{9 \cdot 5 \cdot 9x}=\] Then cancel the 9 from the numerator and the (second) 9 from the denominator:\[\frac{ 5 \cdot \sin 5x }{ 9 \cdot 5x }=\frac{5}{9}\cdot\frac{\sin5x}{5x}\]

OpenStudy (3psilon):

Thank you so much!

OpenStudy (zehanz):

YW!